x<--2 F(x)= {0.25x + 0.5 -25x<2 25x Determine the following: (a) P(X<1.8) (b) P(x>-1.5) (c) P(X<-2) (d) P(-1< X <I)
Answer #1

Solution. Given distribution function Fla) - o - = 0.25 xtos x 2-2 - 2 L x 22 22x The p.dif of o flow= d is, F(x) 0 . 25 -2 c6 P [XY-1.5] = frog dre 21.5 = ² Idre 215 - [ Julis + [2-(-1.5) = 0.8750 . The required probability is 0.8750.© P[XL-2) =P [X L2] - j floul de = 200 – og :. The required probability is o.s. 4 @ Pl_1 < x < 1] = { fcoes doe = 0.5 The req